How many moles of hydrosulphate acid are there in 27.0 g of hydrosulphate acid if the density of hydrosulphate acid is 1.84 g/cubic centimeter. what is the concentration in %m/m of 250 cubic centimeter in weighing 350 g?
1
Expert's answer
2013-03-15T09:42:14-0400
To find the amount of H2SO4 in moles we should use the formula

n(moles) = m(g)/MW(g/mol)
MW(H2SO4) = 2 x 1 + 32 + 16 x 4 = 98 g/mol
n(H2SO4) = 27.0/98.0 = 0.276 mol
The concentration of H2SO4 is
w(%) = m(H2SO4, g)/m(solution of H2SO4, g) x 100%
The weight of 250 cm^3 of solution is
m(H2SO4) = V(cm^3) x p(g/cm^3) = 250 x 1.84 = 460 g
w(%) = 350 / 460 x 100% = 76.1 %

Answer:
n(H2SO4) = 0.276 mol
w(%) = 76.1 %